OpenStudy (anonymous):

Find the regression estimate of the length of a petal that is 1.5 cm wide.

OpenStudy (anonymous):

6B : 1.5-1.8/0.2=-1.2 -1.2*0.25+3.1 =2.8 Ans :2.8